5 x 11=(5x7)+(5xN)
Answer by sparks211a(3)

5 x 11= (5 x 7)+(5 x N)
55=(35)+(5n)
55= 35+5n
-35 -35
20=5n
divide both sides by 5 (5n)
20/5=5n/5
4=n
n=4 is the answer